1. [(e/√x -2√x) - y/√x] = dy/dx
  2. xdy/dx +y=x3y6
  3. dr/dx = tyanchya - r2/coax
  4. y dx/dy = x-yx2siny
1
Expert's answer
2022-01-10T16:41:50-0500

1.

"e^{ -2√x}/\sqrt x - y/\sqrt x = dy/dx"


"y'+\frac{1}{\sqrt{x}}y=e^{ -2√x}/\sqrt x"


"P=\frac{1}{\sqrt{x}},Q=e^{ -2√x}/\sqrt x"


"I.F.=e^{\int Pdx}=e^{\int \frac{1}{\sqrt{x}}dx}=e^{ 2√x}"


"ye^{ 2√x}=\int Q\cdot I.F.dx=\int \frac{e^{- 2√x}}{\sqrt{x}}e^{ 2√x}dx=\int \frac{1}{\sqrt{x}}dx=2\sqrt x+c"


2.

"y'\/y^6+y^{-5}\/x=x^2"

"y^{-5}=v"

"-5y'\/y^6=v'"

"v'-5v\/x=-5x^2"

"IF=e^{-5\int dx/x}=1/x^5"


"v\cdot IF=\int Q\cdot IF dx"


"\frac{1}{(xy)^5}=-5\int dx/x^3"


"\frac{1}{(xy)^5}=\frac{5}{2x^2}+c"


"\frac{1}{x^3y^5}=\frac{5}{2}+c=c_1"


3.

"dr\/dx=-r^2\/cosx"

"-dr\/r^2=dx\/cosx"


"1\/r=ln(tanx+secx)+c"


4.

"dx\/dy-x\/y=-x^2siny"

"u(y)=x(y)\/y,x'=yu'+u"

then:

"y^2u^2siny+yu'=0"

"u'\/u^2=-ysiny"


"-1\/u=ycosy-siny+c"

"-y\/x=ycosy-siny+c"